Out of 1800 students in a school, 750 opted basketball, 800 opted cricket, and the remaining opted table tennis. If a student can opt only one game, find the ratio of
The number of students who opted for basketball to the number of students who opted for table tennis.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Ratio of number of students who opted basketball to the number of students who opted table tennis =
750 / 250 = 3 / 1

Thanks